OpenStudy (anonymous):

HELP!! A parabola is defined by the equation y=-x^2+8x-13 Determine the coordinates of the parabola vertex and draw the parabola. A line l has the equation y=2x-8 Determine the coordinates of each of the intersections between the parabola and the line l

OpenStudy (anonymous):

\[ y=-(x^2 - 8 x) -13\\ y=-(x^2 -8x + 16 -16)-13 y=-(x^2 -8x + 16) +16-13 y=-(x-4)^2+3 Vertex=(4, 3} \]

OpenStudy (anonymous):

@eliassaab where do you get 16 from?

OpenStudy (anonymous):

To complete the square \[ (x-4)^2 = x^2 -8x +16 \] You have to add 16 and take away 16

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

@eliassaab no still extremely confused...